The width of a rectangle is 3 in. less than its length. The perimeter is 24 in. Find the length and width of the rectangle.
Answer by checkley77(12569)

W=L-3
2L+2W=24
2L+2(L-3)=24
2L+2L-6=24
4L=24+6
4L=30
L=30/4
L=7.5 IN. ANS. FOR THE LENGTH.
W=7.5-3=4.5 IN. ANS. FOR THE WIDTH.
PROOF:
2*7.5+2*4.5=24
15+9=24
24=24